OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. x^2 – 36 = 0 and x^2 = 8x – 12 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different?

jhonyy9 (jhonyy9):

for x^2 -36 =0 can using formula a^2 -b^2 =(a-b)(a+b) x^2 -8x +12=0 can be factorizing like (x-6)(x-2)=0 - so than can you continue ?

OpenStudy (anonymous):

no im kinda lost..can you work me through it

jhonyy9 (jhonyy9):

x˘2 -36 =(x-6)(x+6) ok ?

jhonyy9 (jhonyy9):

sorry there is x^2 -36 =(x-6)(x+6)

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so now can you getting the roots of these equations ?

OpenStudy (anonymous):

im confused

jhonyy9 (jhonyy9):

(x-6)(x+6)=0 when this will be satisfied ?

jhonyy9 (jhonyy9):

so when x-6=0 and when x+6=0 yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

but yoou know that x-6=0 so than x=6 and the same when x+6=0 so than x=-6

OpenStudy (anonymous):

okay now what

jhonyy9 (jhonyy9):

youhave there again (x-6)(x-2)=0 can you solve it the same like the first ?

OpenStudy (anonymous):

i got it thanks :)

jhonyy9 (jhonyy9):

ok good luck bye